OpenStudy (anonymous):

Solve for n and select the correct answer. 15n+3/2 = -6 Select one: a. n = 14 b. n = -1 c. n = -3 Incorrect d. n = 19

OpenStudy (zale101):

subtract 3/2 to both sides \[15n+\frac{ 3 }{ 2 } -\frac{ 3 }{ 2}= -6 -\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

i did this on paper but C wasnt the answer

OpenStudy (anonymous):

maybe its B because its supposed to be a negative number right

OpenStudy (anonymous):

am i correct

Nnesha (nnesha):

plug n values into the equation where n is then solve if both side equal then that is your answer

Nnesha (nnesha):

you said B lets check it solve this

Nnesha (nnesha):

are you sure ??? can check your answer choice because that all seem wrong

Nnesha (nnesha):

@prowrestler are you here???

OpenStudy (anonymous):

yes

Nnesha (nnesha):

 did you solve this ???